1. Minutes not distributed. Dean will email them for approval.
2. 2B Weeding at Alexander temporarily halted. Katrina working on E-Resources. Katrina, Suzanne and Jamie working on EAL ongoing projects. Big projects of EAL transfer pretty much done. Suzanne is doing a truck at a time.
MARCIVE and Special Collections (inventory) and e-journal work is being done by Neera. NJ and Special Collections holdings are being consolidated.
Current periodicals have been consolidated to make room for the Café. Students are working on ERIC holdings also. And students are barcoding for the SIP.
SIP is going at LSM and Alexander, Microfilm from the conference room is being addressed, those that were slated for the Alexander MF room have been put out.
Work continues on TechSer and Long overdue books at TSB. Records have been found for most of the books sent over for cataloging. SB and MT will start assisting with Cataloging Department's backlog next.
Microfilm will now be accepted at the Annex. WD of duplicates and Recon is being worked on at the Annex.
Next meeting: [April 13, 2011]: [Pane Room, Alex]Time: [10:00]
Next recorder: [Linda Zuckerman]
Back up recorder: [Katrina Zwaaf]